# Tensor Decomposition Networks for Fast Machine Learning Interatomic Potential Computations

### v1

@@ 6 Summary
In this work, we present Tensor Decomposition Networks (TDNs), a novel framework designed to accelerate the computationally intensive CG tensor product in $\rm{SO}(3)$-equivariant networks through low-rank tensor decomposition. By leveraging CANDECOMP/PARAFAC (CP) decomposition and implementing path-weight sharing, TDNs effectively reduce both parameter count and computational complexity while preserving the expressive power of conventional tensor products. Extensive evaluations on a newly curated PubChemQCR dataset and commonly used OC20 and OC22 benchmarks demonstrate that TDNs achieve comparable predictive accuracy to state-of-the-art models while significantly reducing runtime. The proposed framework provides a plug-and-play alternative to conventional tensor products, making it a promising approach for large-scale molecular simulations and MLIP training.
@@ Acknowledgments
SJ acknowledges support from ARPA-H under grant 1AY1AX000053, National Institutes of Health under grant U01AG070112, and National Science Foundation under grant IIS-2243850. XFQ acknowledges support from the Air Force Office of Scientific Research (AFOSR) under Grant FA9550-24-1-0207. We acknowledge the support of Lambda, Inc. and NVIDIA for providing the computational resources for this project.
@@ Appendix A Multilinear Maps and Tensor Decomposition in Arbitrary Order
Let $N\geq 2$ and let $V_{i}=\mathbb{R}^{d_{i}}$ for $i=1,\dots,N+1$ be finite-dimensional real vector spaces equipped with fixed ordered bases $\{{\bm{e}}_{k}^{(i)}\}_{k=1}^{d_{i}}$. A multilinear map
@@
With respect to the chosen bases, $\widetilde{m}$ is encoded by an $(N\!+\!1)$-way tensor
@@
To lower computational cost, we can approximate ${\bm{M}}$ by a rank-$R$ CP decomposition
@@
where ${\bm{A}}\in\mathbb{R}^{d_{N+1}\times R}$ and ${\bm{B}}^{(i)}\in\mathbb{R}^{d_{i}\times R}\;(i=1,\dots,N)$. Substituting into Eq. 4 and rearranging the summation the we obtain
@@ B.1 Error Bound of CP-Based Tensor Product
{theorem}
Equivariance Error Bound of CP Decompositionequivariance-error Let CG tensor ${\bm{M}}\in\mathbb{R}^{d\times d\times d}$ and $\widehat{{\bm{M}}}$ be the rank-$R$ CP-decomposition-based approximation obtained by Frobenius minimization. For any rotation ${\bm{R}}\in SO(3)$ and any bounded $\rm{SO}(3)$-irreps ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d},\lVert{\bm{x}}\rVert,\lVert{\bm{y}}\rVert\leq C$, we have
[…]
where $R_{T}=\lceil R^{1/3}\rceil$ and $\sigma_{k}^{(n)}$ is the $k$-th singular value of mode-$n$ matricization of ${\bm{M}}$.
@@ B.2 Universality of CP-Based Tensor Product
{theorem}
Universality of CP Decompositionuniversality Consider $\rm{SO}(3)$-irreps ${\bm{x}}\in V_{1}=\mathbb{R}^{d}$, ${\bm{y}}\in V_{2}=\mathbb{R}^{d_{2}}$ and $V_{3}=\mathbb{R}^{d_{3}}$. Fo any $\rm{SO}(3)$-equivariant bilinear map $\Phi$, there exist ${\bm{B}}\in\mathbb{R}^{d_{1}\times R}$, ${\bm{C}}\in\mathbb{R}^{d_{2}\times R}$, ${\bm{A}}\in\mathbb{R}^{R\times d_{3}}$ such that $\Phi$ can be written as
$\Phi({\bm{x}},{\bm{y}})={\bm{A}}({\bm{B}}^{\top}{\bm{x}}\odot{\bm{C}}^{\top}{\bm{y}})\in V_{3},$
if $R$ is sufficient large.
@@
so that $\{{\bm{T}}^{(\ell_{1},\ell_{2},\ell_{3})}_{uvw}\}$ is a basis of $\mathcal{H}$. And then we zero-pad ${\bm{a}}^{(\ell_{3})}_{u},{\bm{b}}^{(\ell_{1})}_{v},{\bm{c}}^{(\ell_{2})}_{w}$ as the vectors in $V_{3}$, $V_{1}$, and $V_{2}$, respectively. Now let
@@ Appendix C Model and Training Configurations
PubChemQCR training.
All experiments on the PubChemQCR benchmarks use a cutoff radius of 4.54.54.5\,4.5Å, the Adam optimizer with initial learning rate $5\times 10^{-4}$, and a ReduceLROnPlateau scheduler with a patience of 2 epochs. Models are trained for up to 100 epochs on PubChemQCR-S and up to 15 epochs on the full PubChemQCR dataset on NVIDIA A100-80GB GPUs.
TDN configuration.
For both PubChemQCR and PubChemQCR-S datasets, TDN employs 6 graph-transformer layers with MLP attention, an irreps-channel dimension of 256, a maximum angular degree $L=1$, and a graph-transformer layer for the force output head. Also, the model is trained with batch size 32 under the above training protocol. For the OC20 IS2RE-Direct task, TDN adopts the same model configuration. For the OC20 S2EF task, TDN uses a similar configuration but with 12 graph-transformer layers and an additional degree-$1$ bond-order orientation (BOO) feature [41] adding to the initial node embedding. For the OC22 IS2RE task, TDN also uses a similar configuration but with 6 graph-transformer layers and an additional degree-$9$ BOO feature [41] adding to the initial node embedding.
@@
Other baselines.
Table 6 summarizes the configurations for all other baseline models. For FAENet we select the “simple” message-passing variant; for MACE we include the residual interaction block. Initial attempts to train Equiformer with its OC20 original settings (6 layers, hidden irreps 256x0e+256x1e or 256x0e+128x1e) failed to converge; we therefore adopt a reduced configuration of 4 layers with irreps 128x0e+64x1e. For all tensor-product-based methods including NequIP [20], MACE [7], PACE [29], Allegro [31], SevenNet [30], and Equiformer [6], only even-parity irreps are retained and $L_{max}=2$ except for Equiformer.
Appendix D Limitations
Our approach accelerates $\rm{SO}(3)$-equivariant tensor products by replacing the full CG tensor product with a low-rank CP decomposition. We substantiate both its theoretical speed-up and its empirical efficacy across multiple benchmarks. A principal limitation is that the minimal rank required to achieve a desired approximation error is not known and determining it exactly is NP-hard. To address this for the CG product, we introduce an empirical rank scheduler; however, this scheduler relies on the CG tensor and must be re-derived for each new equivariant tensor product. In future work, we will investigate the general relationship between group-equivariant tensor products and their optimal rank profiles, with the goal of developing broadly applicable rank-selection strategies.

### v2

@@ 6 Summary
In this work, we present Tensor Decomposition Networks (TDNs), a novel framework designed to accelerate the computationally intensive Clebsch-Gordan (CG) tensor product in $\rm{SO}(3)$-equivariant networks through low-rank tensor decomposition. By leveraging CANDECOMP/PARAFAC (CP) decomposition and implementing path-weight sharing mechanism, TDNs effectively reduce both parameter count and computational complexity while preserving the expressive power of conventional CG tensor products. We also analyze time complexity, derive approximation error bounds, and establish universality of our approach, providing theoretical guarantees. Extensive evaluations on a newly curated PubChemQCR dataset and commonly used OC20 and OC22 benchmarks demonstrate that TDNs achieve comparable predictive accuracy to state-of-the-art models while significantly reducing runtime. The proposed framework provides a plug-and-play alternative to conventional CG tensor products, making it a promising approach for large-scale molecular simulations.
@@ Acknowledgments
SJ acknowledges support from ARPA-H under grant 1AY1AX000053, National Institutes of Health under grant U01AG070112, and National Science Foundation under grant IIS-2243850. XQ acknowledges support from the Air Force Office of Scientific Research (AFOSR) under grant FA9550-24-1-0207. We acknowledge Lambda, Inc. and NVIDIA for providing the computational resources for this project.
@@ Appendix A Multilinear Maps and Tensor Decomposition in Arbitrary Order
Let $N\geq 2$ and let $V_{i}=\mathbb{R}^{d_{i}}$ for $i=1,\dots,N+1$ be finite‑dimensional real vector spaces equipped with fixed ordered bases $\{{\bm{e}}_{k}^{(i)}\}_{k=1}^{d_{i}}$. A multilinear map
@@
With respect to the chosen bases, $\widetilde{m}$ is encoded by an (N+1)(N\!+\!1)‑way tensor
@@
To lower computational cost, we can approximate $\bm{M}$} by a rank‑$R$ CP decomposition
@@
where ${\bm{A}}\in\mathbb{R}^{d_{N+1}\times R}$ and ${\bm{B}}^{(i)}\in\mathbb{R}^{d_{i}\times R}\;(i=1,\dots,N)$. Substituting into Eq.˜4 and rearranging the summation the we obtain
@@ B.1 Error Bound of CP-Based Tensor Product
Theorem B.1 (Equivariance Error Bound of CP Decomposition): Let CG tensor ${\bm{M}}\in\mathbb{R}^{d\times d\times d}$ and $\widehat{{\bm{M}}}$ be the rank-$R$ CP-decomposition-based approximation obtained by Frobenius minimization. For any rotation ${\bm{R}}\in SO(3)$ and any bounded $\rm{SO}(3)$-irreps ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d},\lVert{\bm{x}}\rVert,\lVert{\bm{y}}\rVert\leq C$, we have $\varepsilon({\bm{R}},{\bm{x}},{\bm{y}})\;\leq\;2C^{2}\Bigl(\sum_{n=1}^{3}\sum_{k>R_{T}}\sigma_{k}^{(n)2}\Bigr)^{1/2},$ where $R_{T}=\lceil R^{1/3}\rceil$ and $\sigma_{k}^{(n)}$ is the $k$-th singular value of mode-$n$ matricization of $\bm{M}$}.
@@ B.2 Universality of CP-Based Tensor Product
Theorem B.2 (Universality of CP Decomposition): Consider $\rm{SO}(3)$-irreps ${\bm{x}}\in V_{1}=\mathbb{R}^{d}$, ${\bm{y}}\in V_{2}=\mathbb{R}^{d_{2}}$ and $V_{3}=\mathbb{R}^{d_{3}}$. For any $\rm{SO}(3)$-equivariant bilinear map $\Phi$, there exist ${\bm{B}}\in\mathbb{R}^{d_{1}\times R}$, ${\bm{C}}\in\mathbb{R}^{d_{2}\times R}$, ${\bm{A}}\in\mathbb{R}^{R\times d_{3}}$ such that $\Phi$ can be written as $\Phi({\bm{x}},{\bm{y}})={\bm{A}}({\bm{B}}^{\top}{\bm{x}}\odot{\bm{C}}^{\top}{\bm{y}})\in V_{3},$ if $R$ is sufficiently large.
@@
so that $\{{\bm{T}}^{(\ell_{1},\ell_{2},\ell_{3})}_{uvw}\}$ is a basis of $\mathcal{H}$. And then we zero‐pad ${\bm{a}}^{(\ell_{3})}_{u},{\bm{b}}^{(\ell_{1})}_{v},{\bm{c}}^{(\ell_{2})}_{w}$ as the vectors in $V_{3}$, $V_{1}$, and $V_{2}$, respectively. Now let
@@ Appendix C Model and Training Configurations
PubChemQCR baseline model configuration.
Table˜6 summarizes the configurations for all other baseline models. SchNet [16] and PaiNN [4] are used as implemented in the FAIRChem repository v1. FAENet follows their OC20 release with an $\rm{O}(3)$ stochastic frame and the “simple” message-passing variant. For MACE [8], we include the real-agnostic residual interaction block. For PACE [30], we retain its interaction block and set the edge-booster dimension to 256. For NequIP [24], MACE, Allegro [32], SevenNet [31], and PACE, we adapt the official repositories to PyTorch Geometric and use a Bessel basis with polynomial cutoff smoothing, keeping all numerical settings at their defaults except the irrep settings. For these models, we set identical irrep-channel dimensions according to Table˜6 across irrep blocks. For Equiformer [7], each graph-transformer layer uses 4 attention heads, and the irreps embedding comprises 128 scalars and 64 vectors. All tensor-product-based methods including NequIP, MACE, Allegro, SevenNet, PACE, and Equiformer, only retain even‐parity irreps and use $L_{max}=2$ except Equiformer for fast training.
TDN model configuration.
For PubChemQCR and PubChemQCR-S datasets, TDN employs six graph‑transformer layers with MLP attention, an irreps‑channel dimension of 256, a maximum angular degree $L=1$, and a graph-transformer layer for the force output head. For the OC20 IS2RE-Direct task, TDN adopts the same model configuration and builds the radius graph on the fly with a cutoff of 5.0 and 500 neighbors. For the OC22 IS2RE task, TDN also uses a similar configuration with six graph‑transformer layers, a cutoff of 12.0 with 20 neighbors, and an additional degree-$9$ BOO feature [38] adding to the initial node embedding.
Appendix D Additional Ablation Studies
Ablation study of CP decomposition rank.
To further demonstrate the practical implication of CP-decomposition-based tensor product and the adopted scheduler, we conduct the ablation study of ranks over $n$-body system dataset [41]. In Table˜7 of the n-body system experiment, the TDN is trained with a maximum angular degree $L=2$, three interaction blocks, and a hidden dimension of 72 in 5000 epochs over a single NVIDIA RTX 2080Ti-11GB GPU. As shown in the table, our adopted schedule uses much smaller ranks yet matches the accuracy obtained with the largest ranks, corresponding to the results of rank schedule selection in. Note that the largest rank is deduced from the upper bound in Section˜3.3.
Table 7: TDN Rank‑sweep table of the $n$-body system experiment. The last line is TDN without CP decomposition (CP) and path-weight sharing tensor product (PS).
$R$ $n$-Body Testing MSE Equivariance Error
10 0.0081 0.60
15 0.0064 0.42
20 0.0051 0.21
28 (Our scheduler) 0.0040 0.02
81 (Highest rank) 0.0039 < 0.01
TDN w/o CP + PS 0.0038 -
Time ablation of path-weight sharing and CP decomposition.
Table˜8 presents the time ablation study of path-weight sharing and CP decomposition over the TDN model. As shown in the table, CP decomposition significantly increases the GPU and CPU throughput while path-weight sharing mechanism substantially reduces the parameter count of the model. Note that because TDN removes the depth-wise tensor-product operator, a TDN without CP decomposition and without path-weight sharing in the tensor-product and equivariant-linear layers is not identical to the vanilla Equiformer. All experiments are run on a single NVIDIA A100-80GB GPU and Xeon Gold 6258R processor with a batch size of 128. All experiments are conducted under identical irrep configurations across varying $L$ with 256 irrep-channel dimension, six graph transformer layers, and 8 attention heads of each layer.
Table 8: GPU Throughput and parameter count for Equiformer and TDN variants with or without CP decomposition (CP), path-weight sharing tensor product (PS), path-weight sharing equivariant linear layer (PS-Linear) across maximum degree $L$. Values in parentheses indicate CPU throughput.
$L$ Model / Variant Throughput (samples/sec) Params
1 Equiformer 311.7 (7.5) 12.1
TDN 770.8 (20.2) 4.5
 TDN w/o CP 328.1 4.5
 TDN w/o CP + PS 320.6 5.0
 TDN w/o CP + PS + PS-Linear 317.8 9.1
2 Equiformer 71.9 (2.4) 27.9
TDN 312.4 (8.7) 4.5
 TDN w/o CP 83.7 4.5
 TDN w/o CP + PS 82.5 6.3
 TDN w/o CP + PS + PS-Linear 82.1 14.6
3 Equiformer 26.1 (0.6) 54.7
TDN 220.4 (5.8) 4.5
 TDN w/o CP 26.2 4.5
 TDN w/o CP + PS 25.6 8.9
 TDN w/o CP + PS + PS-Linear 25.6 21.3
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction are detailed in the results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitation in Section˜5.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We present the complete proofs of our theoretical results in Section˜B.1 and Section˜B.2.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide all the details of our method for the reproduction.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We include the code repository link in the abstract.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have specified all the training and test details including data splits, optimization hyperparameters, and model hyperparameters.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: All the datasets we employ are large enough for statistical significance.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We state the computational resources in Appendix˜C.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conform with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We briefly describe the societal impacts in Section˜1.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This is not relevant to this work.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have cited all relevant datasets we used.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We use the existing datasets to train and evaluate our method.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We do not use crowdsourcing in this paper.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We use LLM only for grammar editing.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.